# Topology and Algebra of Bonded Knots and Braids

### v2

@@ 0. Introduction
It is a theme of this paper to compare long bonds, standard bonds and tight bonds in the topological and rigid categories. We present the set of allowed moves for each category and highlight forbidden moves that distinguish bonded links from related concepts like tied links. We also recall and extend invariants for bonded knots: notably, the unplugging technique for the topological vertex category, the tangle insertion for the rigid vertex category, and a Kauffman bracket type polynomial (the bonded bracket) that is invariant under regular isotopy of rigid vertex bonded links. See [24, 28]. Continuing this theme we introduce and study bonded braids as the algebraic counterpart of bonded links. We establish the bonded braid monoid, we point out its relation to the singular braid monoid [4] and we extend it to the bonded braid group. We further establish the interaction between topological bonded links and bonded braids by means of a closure operation, a braiding algorithm and bonded braidoid equivalences. This algebraic structure could be used in encoding the topological structures of bonded links and of objects that they model.
@@
In this paper we consider bonded multi-knotoids in full generality, where the bonds can be knotted and linked, and also up to both equivalence relations: rigid vertex as well as topological vertex equivalence. We also introduce and study bonded braidoids and their interrelation with bonded multi-knotoids.
@@
In regard to interdisciplinary connections, we note that the interaction lines in Feynman diagrams have the formal structure of bonds in our sense. Thus a Feynman diagram can be regarded as a bonded graph (possibly with a knotted embedding in three space). In fact, just such ideas are in back of the work of Kreimer in the book “Knots and Feynman Diagrams” [31] where the bonds in the Feynman diagrams undergo tangle insertion (in our sense) and are thereby associated with specific knots and links. Kreimer suggests that the topological types of these knots and links associated with the diagrams are significantly related to the physical evaluations of the diagrams.
@@
In another direction, we note that current suggestions about knotted glueballs (closed loops of gluon flux related to the structure of protons) can be seen in our context as knotted structures consisting entirely of bonds. The strings of gluon field form highly attracting bonds between quarks in this model [38, 7]. Finally, we point out that there is ongoing research in the interface of molecular biology and the construction of molecules with specified polyhedral shapes [23]. This research also involves bonds to which our modelling applies.
In Section 1, we develop the notion of long bonded links and we establish the allowed isotopy moves in both topological and rigid vertex categories. In Section 2, we introduce standard bonded links, namely, long bonded links with unknotted and unlinked bonds, and we establish that any long bonded link can be isotoped into standard form. We also give a full set of isotopy moves for standard bonded links in both topological and rigid vertex categories. We also identify some forbidden moves due to bonds. In Section 3, we introduce a stricter category of bonded links, called tight bonded links. Tight bonded links are standard bonded links whose bonds do not interact with link arcs. We also present a set of isotopy moves for this setting.
@@
In Section 7 we prove bonded braid equivalence theorems, showing that two bonded braids have topological equivalent closures if and only if they are related by certain moves (Markov moves adapted to bonds) and bonded braid isotopy. We begin with an $L$-move formulation of bonded braid equivalence, by using local cutting and rejoining moves ($L$-moves) on bonded braids. This provides a more fundamental understanding of how bond isotopies translate into more algebraic moves analogous to the classical Markov theorem.
@@
We conclude with a discussion of further directions, including the formulation of a bonded Morse category and the extension of the braidoid–knotoid interaction to the bonded and enhanced bonded settings, which we plan to develop in future work.
Much of this work has been presented in conferences in a number of places. Here is a list of some of these presentations:
SL: (Topological) bonded knots and bonded knotoids; Knots and Proteins Workshop in Ljubljana, Slovenia, 12–14 February 2024 & Quantum Topology Seminar, SUSTech, 10 April 2024.
LK: Topological Bonded Knots and Knotoids; Banff International Research Station, March 21, 2024.
SL: Braids, braidings and braid equivalences; SKCM2 WPI Hiroshima U Winter School, 9–20 December 2024.
SL: Bonded knots, bonded knotoids and bonded DP tangles; SKCM2 Hiroshima U Winter School, 9–20 December 2024.
ID: Bonded Knots and Braids: From Topology to Algebra and Protein Structure; Applied Topology Day, Vrije Universiteit Amsterdam, April 8, 2025.
ID: Bonded knots and bonded braids; Algebraic and Geometric methods of Analysis, AGMA2025, May 26–29, 2025, Ukraine.
@@ 1.1. Definition of Bonded Links
An (oriented) link on $c$ components is an embedding of $c$ (oriented) circles $S^{1}$ in the 3-sphere $S^{3}$. For $c=1$ a link on a single component is a knot. Classical knots and links are considered up to isotopy of the ambient space $S^{3}$, namely orientation-preserving homeomorphisms taking the one knot or link to the other. We usually study knots and links via their diagrams which are regular projections on the plane with over/under conventions at the double points, the crossings, and where isotopy is discretized via the Reidemeister moves (see Figure 5).
@@
A bonded link (or bonded knot, in case of a single component) is a pair $(L,B)$, where $L$ is a link in $S^{3}$, and $B$ is a set of $k$ disjoint arcs properly embedded in the complement $S^{3}\setminus L$, such that each bond arc in $B$ has its two endpoints, called nodes, attached on $L$. The nodes are attached transversely to $L$, with each node attaching at a distinct point on $L$. If $B=\emptyset$, then $(L,\emptyset)$ is just a classical link.
@@ 1.2. Bonded Isotopy and Allowed Moves
Figure illustrates a bonded knot which trivial as topological but non-trivial as rigid bonded knot. It also illustrates a trivial bonded link and a bonded link which is non trivial in either setting.
Clearly, rigid vertex isotopy implies topological vertex isotopy. On the level of diagrams, this means that one can go from a diagram of $(L_{1},B_{1})$ to a diagram of $(L_{2},B_{2})$ by a finite sequence of planar isotopies and a set of local Reidemeister moves involving any types of arcs (link arcs or bonds or both) respecting or not rigid vertex isotopies. To see this, we adapt [24, Theorem 2.1] and [24, Section III] to the bonded link setting and we obtain the following:
@@
    Topological vertex twist moves (TVT), where one end of a bond twists around a strand, as exemplified in Figure 9.
@@
We now present explicitly the moves listed in Proposition 1.5. Planar isotopy moves involving both link arcs and bonds are shown in (Figure 4). Planar isotopy includes also sliding a node along its attaching arc.
@@
Proof of Proposition 1.5. This is an adaptation of [24, Theorem 2.1] to the case of trivalent topological graphs in the particular form of bonded links. The only differences are as follows. Kauffman’s original statement of the basic moves of topological vertex isotopy on embedded graphs includes the following moves (adapted to the bonded category) in place of TVT and VS:
@@
The case of rigid vertex isotopy for bonded links is an adaptation of [24, Section III] to the case of trivalent rigid vertex graphs in the particular form of bonded links and follows analogously. Hence the proof is complete. $\Box$
@@ 1.3. More Diagrammatic Moves for Bonded Links
We now present additional local diagrammatic moves in the setting of topological bonded links, and show that they all follow from the list of basic moves stated in Proposition 1.5 and exemplified in Figures 4567, and 8.
@@
The bonded flype move follows from the TVT move (manifested as bond rotation) and crossing slide moves, as demonstrated in Figure 19. The bonded double flype move is obtained by combining a bonded flype with an R2 move. The bonded double flype move, so also the bonded flype move, follows easily also in the rigid vertex isotopy by the basic moves.
@@ 1.4. Some Restrictions and Forbidden Moves
Another perspective on forbidden moves is given by comparing to the theory of tied links. A tied link is a link together with some ties, which are non-embedded arcs that connect points on the link but can pass freely through the link and even be introduced or removed arbitrarily. Ties were introduced by Aicardi and Juyumaya in [3] as algebraic structures (they correspond to certain algebra elements in a “braid with ties” context). In tied link theory, a tie is like a phantom bond: it does not restrict moves (one can slide a tie through anything). Bonded links differ crucially: bonds are embedded and cannot be created or destroyed by isotopy, nor can they pass through link strands arbitrarily. The forbidden move of Figure 21, for example, would be allowed if the arc were a tie (since a tie can pass through a crossing without issue and even disconnect and reconnect somewhere else), but it is forbidden for a bond. Also, in bonded links one cannot introduce a new bond unless it was already present (whereas ties may be added between any two points, or even multiple ties between the same points). Thus, tied links form a more combinatorially flexible but physically less realistic model, whereas bonded links enforce constraints suitable for modeling fixed interactions (like chemical bonds) that cannot simply appear or vanish under isotopy. For further extensions of tied link theory (including ties in different ambient spaces and braid formulations), see [9, 10].
@@ 2. Standard Bonded Links
Given a bonded link diagram, it is possible to simplify the diagram by applying isotopy moves so that the bonds themselves become unknotted and are presented in a “standard” form. Recall that in a bonded link diagram, bonds may initially be knotted or entangled on their own, forming non-trivial configurations in space. However, by performing appropriate isotopies, one can eliminate such knotting in the bonds and arrange them in a simple configuration where each bond is unknotted. In particular, it is useful to isotope the diagram so that each bond connects two link segments in a simple H-shaped configuration. Consider a small neighborhood around each node on the link: the diagram can be isotoped so that near each bond endpoint the link segment appears roughly vertical and the bond itself attaches like a horizontal rung, together forming a shape reminiscent of the letter “H”, with the bond as the horizontal bar and the link segments as the vertical bars. We refer to such a local configuration as an H-neighborhood of a bond (see Figure 23).
@@
Figure 24. An example.
@@
Two (oriented) standard bonded links are topologically isotopic if and only if any corresponding diagrams of theirs differ by a finite sequence of the moves illustrated in Figures 5 and 25 with all variants.
Similarly, two (oriented) standard bonded links are rigid vertex isotopic if and only if any corresponding diagrams of theirs differ by a finite sequence of the moves illustrated in Figures  5 and 26 with all variants.
Figure 25. Topological isotopy moves between standard bonded links.
@@ 3. Tight Bonded Links
We now introduce a stricter category of bonded links, called tight bonded links. While standard bonded links allow bonds to cross over or under link strands (provided the bonds themselves remain unknotted), tight bonded links forbid such crossings entirely. In a tight bonded link, each bond sits in an H-neighborhood that is free of arcs, as illustrated in Figure 28. We have the following definition:
@@
A tight bonded link diagram is a bonded link diagram in which each bond is presented in standard form and, additionally, no bond crosses any link arc (see Figure 28).
@@
Figure 29. Basic moves for bringing a standard bonded link to a tight form.
@@
Two (oriented) tight bonded links are topologically isotopic in the tight category if and only if any corresponding diagrams of theirs differ by a finite sequence of the moves illustrated in Figures 5 and 30. Similarly, two (oriented) tight bonded links are rigid vertex isotopic in the tight category if and only if any corresponding diagrams of theirs differ by a finite sequence of the moves illustrated in Figures 5 and 31.
Figure 30. Topological isotopy moves between tight bonded links.
Figure 31. Rigid vertex isotopy moves between tight bonded links.
@@
Figure 32. A noded R3 move as a combination of basic moves in the tight category.
@@ 4.1. Topological Bonded Links: Unplugging Operations
For a bonded link $(K,B)$, we view the link together with the bonds as a trivalent embedded graph. At each node (vertex) where a bond attaches, we may perform an unplugging operation: removing one of the edges incident to the vertex (view top row of Figure 33). At a 3-valent vertex there are three such choices. This yields a collection of unknotted (redundant) arcs and classical links corresponding to the various choices of unpluggings. A classical result due to Kauffman [24] asserts, in the language of bonded links:
@@
[…]
@@
The bonded unplugging of a bonded link $(K,B)$ is the unplugging where we simply unplug only the bonds, so we are left with the underlying link $K$.
@@
[…]

### v4

@@ 0. Introduction
It is a theme of this paper to compare long bonds, standard bonds and tight bonds in the topological and rigid categories. We present the set of allowed moves for each category and highlight forbidden moves that distinguish bonded links from related concepts like tied links. We also recall and extend invariants for bonded knots: notably, the unplugging technique for the topological vertex category, the tangle insertion for the rigid vertex category, and a Kauffman bracket type polynomial (the bonded bracket) that is invariant under regular isotopy of rigid vertex bonded links. See [23, 26].
Continuing this theme we introduce and study bonded braids as the algebraic counterpart of bonded links. We establish the bonded braid monoid, we point out its relation to the singular braid monoid [4] and we extend it to the bonded braid group. We further establish the interaction between oriented topological bonded links and bonded braids by means of a closure operation, a bonded braiding algorithm (bonded analogue of the classical Alexander theorem) and bonded $L$-move equivalences as bonded analogues of the classical Markov theorem. The classical Alexander and Markov theorems [2, 32] relate knots and links to braids and the Artin braid groups, and translate their isotopy to an algebraic equivalence among braids. So, the algebraic structure of braids together with the two theorems above furnish the necessary basis for encoding knotted objects by words in braid generators and for the potential construction of knot invariants using algebraic tools. A great paradigm of this approach is the breakthrough work of V.F.R. Jones in the construction of the Jones polynomial (see[21] and references therein), the ambient isotopy equivalent to the Kauffman bracket polynomial. The $L$-move equivalent formulation [30] of the Markov theorem provides a geometric as well as algebraic approach to the classical braid equivalence.
@@
In this paper we focus on bonded knots/knotoids and bonded braids/braidoids. Our motivation for exploring the diagrammatic settings and equivalences of bonded knots/knotoids comes mainly from the remarkable applications so far and their interest as mathematical objects. Then, a new diagrammatic setting leads naturally to the fundamental questions about existence or not of related braid structures. The algebraic structure of bonded braids can be used for encoding the objects that they model with words in the generators, after applying a braiding algorithm for turning them isotopically into closed bonded braids. Topological vertex isotopy translates into a bonded braid equivalence generated by moves between bonded braids. One can exploit algebraic tools for constructing topological invariants in order to distinguish our topological objects. The $L$-moves generating the classical braid equivalence are fundamental, so that they provide an adaptive frame for formulating braid equivalences in other diagrammatic settings. For ensuring a sufficient set of moves for generating a braid equivalence one has to examine all algorithmic choices and moves in the diagrammatic setting, and finding all of them can be very subtle. See for example [30, 29, 24] for different diagrammatic settings. This is also the case here for pinning down the bonded $L$-moves, which augment the classical $L$-moves for generating the bonded braid equivalence. We further describe an algorithm for turning a bonded knotoid into a bonded braidoid and we also extend the concept of bond by introducing two types of mutually cancelling bonds, the enhanced bonds. The definition and study of bonded braids/braidoids and their relation to bonded knots/knotoids form the core of this work.
@@
In regard to interdisciplinary connections, we note that the interaction lines in Feynman diagrams have the formal structure of bonds in our sense. Thus a Feynman diagram can be regarded as a bonded graph (possibly with a knotted embedding in three-space). In fact, just such ideas are in back of the work of Kreimer in the book “Knots and Feynman Diagrams” [28] where the bonds in the Feynman diagrams undergo tangle insertion (in our sense) and are thereby associated with specific knots and links. Kreimer suggests that the topological types of these knots and links associated with the diagrams are significantly related to the physical evaluations of the diagrams.
@@
In another direction, we note that current suggestions about knotted glueballs (closed loops of gluon flux related to the structure of protons) can be seen in our context as knotted structures consisting entirely of bonds. The strings of gluon field form highly attracting bonds between quarks in this model [35, 6]. Finally, we point out that there is ongoing research in the interface of molecular biology and the construction of molecules with specified polyhedral shapes [22]. This research also involves bonds to which our modelling applies. The paper is structured as follows. In Section 1, we develop the notion of long bonded links and we establish the allowed isotopy moves in both topological and rigid vertex categories. In Section 2, we introduce standard bonded links, namely, long bonded links with unknotted and unlinked bonds, and we establish that any long bonded link can be isotoped into standard form. We also give a full set of isotopy moves for standard bonded links in both topological and rigid vertex categories. We also identify some forbidden moves due to bonds. In Section 3, we introduce a stricter category of bonded links, called tight bonded links. Tight bonded links are standard bonded links whose bonds do not interact with link arcs. We also present a set of isotopy moves for this setting.
@@
In Section 7, we prove bonded braid equivalence theorems, showing that two bonded braids have topologically equivalent closures if and only if they are related by certain moves adapted to bonds and (tight) bonded braid isotopy. We begin with the adaptation of classical $L$-moves to (tight) bonded braids, which we further extend to the more subtle (tight) bonded $L$-moves. These moves provide a more fundamental understanding of how bonded isotopies translate into more algebraic moves analogous to the classical Markov theorem.
@@
We conclude with a discussion of further directions, including the further algebraic exploration of the bonded braid equivalences, the extension of the study to the plat closure of bonded braids, the formulation of a bonded Morse category and the extension of the braidoid–knotoid interaction to the bonded and enhanced bonded settings, which we plan to develop in future works.
Much of this work has been presented in conferences in a number of places, such as University of Ljubljana, Banff International Research Station, SKCM2 WPI Hiroshima University, Vrije Universiteit Amsterdam, Odessa National University of Technology.
@@ 1. Bonded Links
Knotted objects with bonds and their isotopy moves have been introduced for the first time in the context of bonded knotoids [15] and then as bonded knots with distinguishable bonds in [13]. In this section we consider bonded knots and links and their isotopy moves in full generality, in the sense that the bonds can be knotted and linked.
@@ 1.1. Definition of Bonded Links
An (oriented) link on $c$ components is an embedding of $c$ (oriented) circles $S^{1}$ in the 3-sphere $S^{3}$. For $c=1$ a link on a single component is a knot. Classical knots and links are considered up to isotopy of the ambient space $S^{3}$, namely orientation-preserving homeomorphisms taking the one knot or link to the other. We usually study knots and links via their diagrams which are regular projections on the plane with over/under conventions at the double points, the crossings, and where isotopy is translated into planar isotopy (see left hand illustration of Figure 4) and the Reidemeister moves (see Figure 5).
@@
A bonded link (or bonded knot, in case of a single component) is a pair $(L,B)$, where $L$ is a link in $S^{3}$, and $B$ is a set of $k$ disjoint arcs properly embedded in the complement $S^{3}\setminus L$, such that each bond arc in $B$ has its two endpoints, called nodes, attached on $L$. The nodes are attached transversely to $L$, with each node attaching at a distinct point on $L$. If $B=\emptyset$, then $(L,\emptyset)$ is just a classical link. In this general setting, where bonds can be knotted and linked, they will be referred to as long bonds.
@@ 1.2. Bonded Isotopy and Allowed Moves
It is easy to create examples of bonded knots which are trivial as topological but non-trivial as rigid bonded knots. Clearly, rigid vertex isotopy implies topological vertex isotopy. On the level of diagrams, this means that one can go from a diagram of $(L_{1},B_{1})$ to a diagram of $(L_{2},B_{2})$ by a finite sequence of planar isotopies and a set of local Reidemeister moves involving any types of arcs (link arcs or bonds or both) respecting or not rigid vertex isotopies. To see this, we adapt [23, Theorem 2.1] and [23, Section III] to the bonded link setting and we obtain the following:
@@
    Topological vertex twist moves (TVT), as exemplified in Figure 9.
@@
We now present explicitly the moves listed in Proposition 1.5. Planar isotopy moves involving both link arcs and bonds are shown in Figure 4. Planar isotopy includes also sliding a node along its attaching arc.
@@
Proof of Proposition 1.5. This is an adaptation of [23, Theorem 2.1] to the case of trivalent topological graphs in the particular form of bonded links. The only differences are as follows. Kauffman’s original statement of the basic moves of topological vertex isotopy on embedded graphs includes the following moves (adapted to the bonded category) in place of TVT and VS, see Figure 11:
@@
The case of rigid vertex isotopy for bonded links is an adaptation of [23, Section III] to the case of trivalent rigid vertex graphs in the particular form of bonded links and follows analogously. Hence the proof is complete. $\Box$
@@ 1.3. More Diagrammatic Moves for Bonded Links
We now present additional local diagrammatic moves in the setting of topological bonded links, and show that they all follow from the list of basic moves stated in Proposition 1.5 and exemplified in Figures 49.
@@
The bonded flype move follows from the TVT move (manifested as bond rotation, that is, mTVT moves) and crossing slide moves, as demonstrated in Figure 19. The bonded double flype move is obtained by combining a bonded flype with an R2 move. The bonded double flype move, so also the bonded flype move, follows easily also in the rigid vertex isotopy by the basic moves.
@@ 1.4. Some Restrictions and Forbidden Moves
Another related diagrammatic setting is the theory of tied links [3, 10], links equipped with ties. A tie is a non-embedded simple arc connecting different points of a link, whose ends can slide along the arcs they are attached but remain on the link. Ties are like phantom bonds: they can pass freely through one another and through the arcs of the link. Two ties can even merge into one if they connect the same arcs. Bonded links differ crucially: bonds are embedded arcs and cannot be created or destroyed by isotopy, nor can they pass through link arcs or through themselves arbitrarily.
@@ 2. Standard Bonded Links
Given a bonded link diagram, it is possible to simplify the diagram by applying isotopy moves so that the bonds themselves become unknotted and are presented in a “standard” form. Knotted objects with bonds in standard form and their isotopy moves have been first introduced in the theory of bonded knotoids [15] and later as bonded knots with distinguishable bonds in [13].
Recall that in a bonded link diagram, bonds may initially be knotted or entangled on their own, forming non-trivial configurations in space. However, by performing appropriate isotopies, one can eliminate such knotting in the bonds and arrange them in a simple configuration where each bond is unknotted. In particular, it is useful to isotope the diagram so that each bond connects two link segments in a simple H-shaped configuration. See Figure 23 for an example. Consider a small neighborhood around each node on the link: the diagram can be isotoped so that near each bond endpoint the link segment appears roughly vertical and the bond itself attaches like a horizontal rung, together forming a shape reminiscent of the letter “H”, with the bond as the horizontal bar and the link segments as the vertical bars. We refer to such a local configuration as an H-neighborhood of a bond (see Figure 23).
@@
Figure 24. An example of bond contraction.
@@
Two (oriented) standard bonded links are topologically isotopic if and only if any corresponding diagrams of theirs differ by a finite sequence of the moves illustrated in Figures 4, 5 and 25 with all variants.
Similarly, two (oriented) standard bonded links are rigid vertex isotopic if and only if any corresponding diagrams of theirs differ by a finite sequence of the moves illustrated in Figures 4, 5 and 26 with all variants.
Figure 25. Topological vertex isotopy moves between standard bonded links.
@@ 3. Tight Bonded Links
We now consider a stricter category of bonded links, called tight bonded links. While standard bonded links allow bonds to cross over or under link strands (provided the bonds themselves remain unknotted), tight bonded links forbid such crossings entirely. See Figure 28 for an exmaple. In a tight bonded link, each bond sits in an H-neighborhood that is free of arcs, as illustrated in Figure 28. We have the following definition (compare with [15] and [13]):
@@
A tight bonded link diagram is a bonded link diagram in which each bond is presented in standard form and, additionally, no bond crosses any link arc.
@@
Figure 29. A standard bond can be brought to tight form by vertex slide moves.
@@
Two (oriented) tight bonded links are topological vertex isotopic in the tight category if and only if any corresponding diagrams of theirs differ by a finite sequence of planar isotopies and the Reidemeister moves for the link arcs (recall Figures 4, 5), and the moves illustrated in Figure 30. Similarly, two (oriented) tight bonded links are rigid vertex isotopic in the tight category if and only if any corresponding diagrams of theirs differ by a finite sequence of planar isotopies and the Reidemeister moves for the link arcs, and the moves illustrated in Figure 31.
Figure 30. Basic topological isotopy moves between tight bonded links.
Figure 31. Basic rigid vertex isotopy moves between tight bonded links.
@@
Figure 32. An external arc slide move as a combination of basic moves in the tight category.
@@ 4.1. Topological Bonded Links: Unplugging Operations
For a bonded link $(K,B)$, we view the link together with the bonds as a trivalent embedded graph. At each node (vertex) where a bond attaches, we may perform an unplugging operation: removing one of the edges incident to the vertex as shown in the top row of Figure 33. At a 3-valent vertex there are three such choices. This yields a collection of unknotted (redundant) arcs and classical links corresponding to the various choices of unpluggings. A classical result due to Kauffman [23] asserts, in the language of bonded links:
@@
  • Consequently, any ambient isotopy invariant applied to $\mathcal{L}_{(K,B)}$ is also a topological vertex isotopy invariant of $(K,B)$.
@@
The bonded unplugging of a bonded link $(K,B)$ is the unplugging where we unplug only the bonds, so we are left with the underlying link $K$.
@@
The underlying link $K$ of a bonded link $(K,B)$ is a topological vertex isotopy invariant of $(K,B)$.